Solve
2y^2 - 3y - 6 = 0

Answer by Nate(3500)

2y^2 - 3y - 6 = 0
2y^2 - 3y = 6
y^2 - 3y/2 = 3
(y - 3/4)^2 = 3 + 9/16 = 48/16 + 9/16 = 57/16
y = [3 +- sqrt(57)]/4